Alice's Restaurant has a total of 205 seats. The number of seats in the non-smoking section is 73 more than twice the number of seats in the smoking section. How many seats are in each section?
Answer by algebrahouse.com(1659)

"Alice's Restaurant has a total of 205 seats. The number of seats in the non-smoking section is 73 more than twice the number of seats in the smoking section. How many seats are in each section?"

x = smoking seats
2x + 73 = non-smoking seats {73 more than twice the smoking seats}

x + 2x + 73 = 205 {smoking + non-smoking = total seats}
3x + 73 = 205 {combined like terms}
3x = 132 {subtracted 73 from both sides}
x = 44 {divided both sides by 3}
2x + 73 = 161 {substituted 44, in for x, into 2x + 73}

44 smoking seats
161 non-smoking seats

For more help from me, visit: www.algebrahouse.com